In divisional income statements prepared for Demopolis Company, the Payroll Department costs are charged back to user divisions on the basis of the number of payroll distributions, and the Purchasing Department costs are charged back on the basis of the number of purchase requisitions. The Payroll Department had expenses of $33,952, and the Purchasing Department had expenses of $12,470 for the year. The following annual data for Residential, Commercial, and Government Contract divisions were obtained from corporate records:


Residential

Commercial
Government
Contract
Sales $ 261,000 $ 345,000 $ 793,000
Number of employees:
Weekly payroll (52 weeks per year) 140 80 85
Monthly payroll 28 39 26
Number of purchase
requisitions per year 1,800 1,300 1,200

a. Determine the total amount of payroll checks and purchase requisitions processed per year by the company and each division.

Residential Commercial Government Contract Total
Number of payroll checks:
Weekly payroll
Monthly payroll
Total
Number of purchase requisitions per year

b. Using the activity base information in (a), determine the annual amount of payroll and purchasing costs charged back to the Residential, Commercial, and Government Contract divisions from payroll and purchasing services. Do not round your interim calculations, round your answers to two decimal places, if required.

Service department charge rates:
Payroll Department $ per payroll check
Purchasing Department $ per purchase requisition


Residential Commercial Government Contract Total
Service department charges:
Payroll Department $ $ $ $
Purchasing Department
Total $ $
Answer #1

SOLUTION

(A)

Residential Commercial Government Contract Total
Number of payroll checks:
Weekly payroll (140*52) = 7,280 (80*52) = 4,160 (85*52) = 4,420 15,860
Monthly payroll (28*12) = 336 (39*12) = 468 (26*12) = 312 1,116
Total 7,616 4,628 4,732 16,976
Number of purchase requisitions per year 1,800 1,300 1,200 4,300

(B)

Service department charge rates:
Payroll Department (33,952/16,976) $2 per payroll check
Purchasing Department (12,470/4,300) $2.9 per purchase requisition
Residential Commercial Government Contract Total
Service department charges:
Payroll Department (7,616*2) = 15,232 (4,628*2) = 9,256 (4,732*2) = 9,464 33,952
Purchasing Department (1,800*2.9) = 5,220 (1,300*2.9) = 3,770 (1,200*2.9) = 3,480 12,470
Total 20,452 13,026 12,944 46,422